OpenStudy (anonymous):

Help me turn this term into simplest form: \[(8a^-3)^\dfrac{-2}{3}\] \[(\dfrac{8}{a^3})^\dfrac{-2}{3}\]

OpenStudy (unklerhaukus):

\[\left(\frac{x}{y}\right)^m=\frac{x^m}{y^{m}}\]

OpenStudy (anonymous):

\[\frac{8^\frac{-2}{3}}{a^{3}}\] Um, wait how would I do it for the denominator? \[3*\frac{-2}{3} = -2\] \[\frac{8^\frac{-2}{3}}{a^{-2}}?\]

OpenStudy (unklerhaukus):

that's right \[(8/a^3)^{−2/3}=\frac{8^{-2/3}}{a^{3\times{-2/3}}}=\frac{8^{-2/3}}{a^{-2}}\] now to simplify the numerator , can you change 8 into something^3?

OpenStudy (anonymous):

\[2^3 = 8\]

OpenStudy (unklerhaukus):

yeah ,

OpenStudy (anonymous):

Wait! I get it. 3 * -2/3 = -2 So: \[\frac{2^{-2}}{a^{-2}}\]

OpenStudy (unklerhaukus):

very good

OpenStudy (unklerhaukus):

now you can simplify further \[\frac{x^n}{y^n}=\left(\frac{x}{x}\right)^n\] and \(\left(\dfrac{x}{y}\right)^{-p}=\left(\dfrac{y}{x}\right)^p\)

OpenStudy (anonymous):

\[\frac{2^{-2}}{a^{-2}}\Rightarrow \frac{a^2}{4}?\]

OpenStudy (unklerhaukus):

*\[\frac{x^n}{y^n}=\left(\frac{x}{y}\right)^n\]

OpenStudy (unklerhaukus):

That is right you've got it!

OpenStudy (anonymous):

Thank you for teaching me :-).